Sal Amato operates a residential landscaping business in an affluent suburb of St. Louis. In an effort to provide quality service, he has concentrated solely on the design and installation of upscale landscaping plans (e.g., trees, shrubs, fountains, and lighting). With his clients continually requesting additional services, Sal recently expanded into lawn maintenance, including fertilization.

The following data relate to his first year’s experience with 55 fertilization clients:

  • Each client required six applications throughout the year and was billed $40.00 per application.

  • Two applications involved Type I fertilizer, which contains a special ingredient for weed control. The remaining four applications involved Type II fertilizer.

  • Sal purchased 5,600 pounds of Type I fertilizer at $0.59 per pound and 10,600 pounds of Type II fertilizer at $0.46 per pound. Actual usage amounted to 3,800 pounds of Type I and 8,100 pounds of Type II.

  • A new, part-time employee was hired to spread the fertilizer. Sal had to pay premium wages of $12.10 per hour because of a very tight labor market; the employee logged a total of 177 hours at client residences.

  • Based on previous knowledge of the operation, articles in trade journals, and conversations with other landscapers, Sal established the following standards:

  • Fertilizer purchase price per pound: Type I, $0.56; Type II, $0.48
  • Fertilizer usage: 46 pounds per application
  • Typical hourly wage rate of landscape personnel: $9.60
  • Labor time per application: 40 minutes
  • The operation did not go as smoothly as planned, with customer complaints actually much higher than expected.

Required:

  1. 1. Compute Sal’s direct-material variances for each type of fertilizer.

  2. 2. Compute the direct-labor variances.

  3. 3-a. Compute the actual cost of the client applications. (Note: Exclude any fertilizer in inventory, as remaining fertilizer can be used next year.)

  4. 3-b. Calculate the profit or loss of Sal’s new lawn fertilization service.
Answer #1

Ans:-

Part 1)

The direct material-variances for each type of fertilizer are calculated as below:

Direct Material Price Variance

Direct Material Price Variance = Actual Quantity Purchased*(Actual Price - Standard Price)

Direct Material Price Variance (Fertilizer I) = 5,600*(.59- .56) = $168 (Unfavorable)

Direct Material Price Variance (Fertilizer II) = 10,600*(.46 - .48) = $212 (Favorable)

_____

Direct Material Quantity Variance

DIrect Material Quantity Variance = Standard Price*(Actual Quantity Purchased - Standard Quantity Purchased for Actual Output)

Direct Material Quantity Variance (Fertilizer I) = .56*(3,800 - 55*2*46) = $630 (Favorable)

Direct Material Quantity Variance (Fertilizer II) = .48*(8,100 - 55*4*46) = $969.60 (Favorable)

______

Part 2)

The direct labor variances are computed as below:

Direct Labor Rate Variance = Actual Hours Used*(Actual Rate Per Hour - Standard Rate Per Hour) = 177*(12.10- 9.60) = $422.50 (Unfavorable)

Direct Labor Efficiency Variance = Standard Rate Per Hour*(Actual Hours Used - Standard Hours Used for Actual Production) = 9.60*(177 - 55*6*40/60) = $412.80 (Favorable)

______

Part 3-a)

The actual cost of applications is determined as follows:

Actual Cost of Type I Fertilizer (3,900*.59) 2,301.00
Actual Cost of Type II Fertilizer 8,100*.46) 3,726.00
Direct Labor Cost (177*12.10) 2,141.70
Total Actual Cost of Client Applications $8,168.70

______

Part 3-b)

The profit or loss Sal's new lawn fertilization service is calculated as below:

Total Revenue (6*55*40) 13,200.00
Total Actual Cost of Client Applications 8,168.70
Profit of Sal's New Lawn Fertilization Service $5,031.30